If we have a linear equation what is the power of the variable(s) involved

1
Expert's answer
2022-03-03T17:33:31-0500

The linear equation is an equation that may be put in the form

"a_1x_1+a_2x_2+\ldots+a_nx_n + c = 0," where "a_1,\ldots, a_n, c" are constants.

So every variable "x_1, \ldots, x_n" should be in power less or equal to 1. Every involved in the equation variable should be in power 1.

For example, "2x + \sqrt7\cdot y - 1.13 z+ 6 = 0" is a linear equation, but "2x^2 - 1= 0" is not linear.